A rectangular beam 10m long is simply supported at its ends and is acted upon by a 10kN/m uniformly distributed load throughout its length. Find the size of the beam if the thickness is twice that of its width. Take the allowable bending stress of the material as 40MN/m
1
Expert's answer
2018-08-14T10:14:08-0400
Maximum bending momtnt
M=(wt^2)/8
M=(10×〖10〗^2)/8=125 kNm
Then
M=fz=f×1/6 bd^2
d=2b
M=f×4/6 b^3

b^3=6M/4f
b^3=(6×125×1000)/(4×40×〖10〗^6 )
b=0.167 m
d=2×0.167=0.335 m